# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

### v2

@@ IV-C Guided DDPM Fine-Tuning
Figure 4: Illustration of SSAT performance using our LCS-KM method on image benchmarks with varying boundary and unlabeled data reduction ratio parameters with respect to: (a) and (c) DDPM-generated CIFAR-10 data with $\alpha=20\%$, (b) and (d) 531531K extra SVHN data with $\alpha=10\%$. For each figure, we vary the considered ratio parameter while keeping all the remaining hyperparameters fixed.
V Additional Analyses
In this section, we present additional analyses to provide further insights into our methods. We first evaluate their performance on a real-world medical application dataset (Section V-A), and then visualize the selected data distribution in the latent space to qualitatively understand their ability in identifying critical data points (Section V-B). Finally, we provide ablation studies on key hyperparameters of our methods. Due to page limit, additional experiments on other hyperparameters, generalizability under different training schemes and perturbations, and the effect of intermediate model quality in low-label regimes, are provided in Appendix D.
V-A Application to Real-World Medical Data
To test the generalization of our findings, we extend our study to a real-world healthcare application. COVIDGR [30] is used as the labeled dataset, and CoronaHack [31] serves as the source of the unlabeled data. Specifically, COVIDGR contains $852$ chest X-ray images, evenly distributed across two classes: $426$ COVID-19 positive (RT-PCR confirmed) and $426$ negative cases. It is then split into $80\%$ training and $20\%$ test sets. CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios. Further details of the experimental setup is discussed in Appendix B.
The results are demonstrated in Figure 2, which confirms the effectiveness of our selection scheme in enhancing robustness while optimizing efficiency. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b). In comparison, using a smaller unlabeled dataset leads to faster convergence, underscoring the benefits of early stopping. Specifically, the best performance of $56\%$ is reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM (Figure 2b). Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve the peak robust accuracy of $58\%$ (Figure 2d). These findings confirm that using a smaller but strategically selected unlabeled data subset can achieve robustness comparable to that of the whole dataset while significantly reducing computational costs. These results highlight the practical relevance of our data reduction scheme, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
V-B Comparative Visualizations of Our Methods
To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10. We first extract their latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model. For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships. Specifically, we first perform Principal Component Analysis (PCA) [32]. By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data. In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [33] to project the $50$-dimensional PCA-reduced features onto a two-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points. After projecting onto the two-dimensional space, we visualize the set of unlabeled data returned by our selection schemes under $\alpha=10\%$.
Figure 3 presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy. Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT. In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes. Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers that offer limited benefit for refining the decision boundary. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a).
In comparison, LCS-GMM and LCS-KM yield more structured sampling patterns. While LCS-GMM can roughly capture regions where classes overlap, it assumes that each cluster follows approximately Gaussian distributions, an assumption that may fail in real-world settings. As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b). A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical examples for training. In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM.
Finally, our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner. By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c). Compared with LCS-GMM, a notable pattern is that the unlabeled data picked by LCS-KM are aligned more linearly. This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer. We hypothesize that such a local linear property potentially explains why our latent clustering-based approach with k-means always achieves the best performance compared to the other two alternatives, which is worth further investigation.
V-C Hyperparameter Sensitivity
The success of our methods relies on a few key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$. We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally. Due to page limits, the detailed settings of these experiments and the ablations on $\gamma$ are deferred to Appendix C-A.
Ratio of Boundary Data $\bm{\beta}$. Concentrating exclusively on data points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away, where the ratio is determined by $\beta$. Figure 4a illustrates how robust accuracy varies with $\beta$ on CIFAR-10 using $20\%$ generated data from DDPM. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure 4b illustrates how robust accuracy varies with $\beta$ on SVHN using $10\%$ extra data from the same dataset. Due to a smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. As a result, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, using a smaller $\beta$ usually leads to better robust accuracy. Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.
@@ V-C Hyperparameter Sensitivity
Data Reduction Ratio $\bm{\alpha}$. Figures 4c and 4d describe how accuracies vary with the ratio parameter $\alpha$. We analyze how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary. When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
(a) Fine-Tuning Epochs $S$
(b) Regularization Strength $\lambda$
Figure 5: Clean and robust accuracy curves of SSAT with LCG-KM using generated data on CIFAR-10 under $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
Guided DDPM Finetuning. We conduct ablation studies on our guided DDPM fine-tuning approach to better illustrate its behavior. In particular, we focus on two key hyperparameters, the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$, and study their impact on accuracy and robustness performance when training SSAT algorithms on CIFAR-10 with LCG-KM and $\alpha=10\%$. We also investigate the quality of the data samples generated by our fine-tuned DDPM using both quantitative and qualitative metrics.
Figure 5a illustrates our ablation results on $S$. We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process. Figure 6 visualizes the images generated by our DDPM models with different fine-tuning epochs. We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained for CIFAR-10. Training for too many epochs can lead to a degradation in image quality and loss of diversity, thereby adversely influencing the final robust accuracies. In addition, Figure 5b shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$. Our findings show that choosing $\lambda$ within the range of $[0.4,0.6]$ results in the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
Figure 6: Visualization of generated data across different fine-tuning epochs
VI Conclusion and Future Work
In this work, we proposed various unlabeled data reduction schemes to improve the efficiency of SSAT while preserving robust accuracy. In particular, we emphasized that strategically selecting or generating more boundary-adjacent data points can significantly improve the efficiency of semi-supervised adversarial training algorithms, without losing their robustness advantages. Among our proposed methods, we found that latent-based approaches using k-means clustering, including LCS-KM and LCG-KM, perform the best, as they reduce computational costs and memory requirements.
Limitation & Future Work. We acknowledge that our work has several limitations. While the selection scheme with LCS-KM consistently achieves the best performance, its success requires carefully choosing a few hyperparameters (e.g., to avoid overfitting). Future studies can leverage AutoML algorithms to automatically determine these hyperparameters. Additionally, it remains unclear whether our best-performing latent-clustering-based methods (i.e., LCS-KM and LCG-KM) can be further optimized to achieve better performance. Theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful. Furthermore, our guided DDPM fine-tuning approach proposes to adapt a pre-trained model for boundary-seeking data generation. An interesting future direction is to investigate whether training-free sampling-based guidance [34] or more advanced generative models [35, 36] can enhance the efficiency and robustness of SSAT. Despite these limitations, we believe our work can serve as an important initial step towards the more efficient and scalable development of robust models.
Availability
The implementation of the latent-based data reduction methods and all our experiments are available as open-source code at: https://github.com/TrustMLRG/EfficientSSAT.
Appendix A Related Work
A-A Adversarially Robsut Learning
Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [1]. The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications. Earlier-proposed attacks like fast gradient sign method (FGSM) [2] and projected gradient descent (PGD) [7] introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks [37], which search for the minimal perturbation that causes misclassification. Recent attacks like Auto-PGD [38] automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack [17] is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
Adversarial Training. On the defense side, adversarial training and its variants are the most popular methods for achieving robustness. For instance, Madry et al. proposed to train DNNs on adversarially perturbed inputs using PGD attacks [7], while Zhang et al. introduced TRADES [9], which balances the trade-off between robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications. In particular, Rice et al. illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training [28], where early stopping is shown to be effective in alleviating such an issue. Techniques, such as Free AT [18] and Fast AT [19], offer promising solutions to reduce the computation for adversarial training. Nevertheless, they suffer from decreased robustness performance.
Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [16]. As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data. In a similar vein, Hua et al. advocated for a more targeted approach [24], recommending that PGD training be applied primarily to examples situated near the decision boundaries. This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks. Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
Self-Supervised Adversarial Training. Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research [39]. In the adversarial context, Schmidt et al. [10] highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning. They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly. Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training [11, 40, 41, 12]. In particular, Carmon et al. proposed robust self-training methods to address the sample complexity issue [11], which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework incorporating labeled and unlabeled data [41], while providing formal guarantees on adversarial robustness.
Leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature [13, 14, 15]. For instance, Gowal et al. employed various unconditional generative models [13], including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further. While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
A-B Data-Efficient Deep Learning
Dataset Distillation. Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points. For instance, Wang et al. proposed a foundational approach to dataset distillation [42], proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching [43], where distilled data points are optimized to match the gradients of the original dataset. This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data. Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [44], facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. On the other hand, coreset selection [45, 46] aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset. These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts [20], while Xia et al. introduced the concept of “moderate coreset” [21], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [22].
In the adversarial context, Dolatabadi et al. proposed adversarial coreset selection [47], offering a task-specific solution to enhance the computational efficiency of adversarial training. In particular, the method constructs a compact data subset by minimizing the gradient approximation error. The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals. While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges. These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time. In such settings, repeatedly selecting data during training may become computationally infeasible. In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately 7373K images and an extra training set of around 531531K images. Initially, the model is trained on the 7373K labeled images. To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model [13].
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture [48] in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. [9], with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. [11] except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with 11M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al. [11] for better comparison. We use step size $0.005$, number of attack steps $K=100$. We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in 8080M-TI do not correspond to CIFAR-10 image categories. Carmon et al. [11] used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model with 5050K labeled data and use this model to select data from the 500500K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments utilizing 11M synthetical data generated by the DDPM model [13]. For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [49] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section D-A when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. [11] except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. [11] for fair comparisons. We use step size $\alpha=0.01$, number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
Medical Application. Below, we detail the setup employed for our experiments on the medical datasets outlined in Section V-A. For this experiment, we trained models using a ResNet-18 [50] architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained with SSAT using the Adam optimizer [51] with a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Appendix C Additional Experiments
@@
Figure 7: Illustration of clean and robust accuracy curves of SSAT algorithms with LCS-KM by varying per-batch ratio $\gamma$, which defines how much labeled and unlabeled data are included in each mini-batch, on: (a) CIFAR-10 with $20\%$ extra unlabeled data selected from 11M generated images using DDPM, and (b) SVHN with $10\%$ selected from the external 531531K SVHN images.
TABLE IV: Comparison results of semi-supervised adversarial training algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
$\textbf{Ratio }\bm{\alpha}$ Method Clean Accuracy ($\%$) Robust Accuracy ($\%$) Number of Epochs

### v3

@@ Appendix C Additional Experiments
Figure 6: Figures (a) and (b): Illustration of SSAT performance using LCS-KM by varying unlabeled data ratios $\alpha$ for CIFAR-10 with DDPM generation ($\beta=0.4$) and SVHN with external data ($\beta=0.6$). Figures (c) and (d): Clean and robust accuracy curves of SSAT with LCS-KM by varying per-batch ratio $\gamma$ on CIFAR-10 with 11M generated images using DDPM ($\alpha=20\%$, $\beta=0.4$) and on SVHN with external 531531K images ($\alpha=10\%$, $\beta=0.6$).
C-A Other Ablations on Hyperparameters
In Section VI-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. We vary one hyperparameter at a time while keeping the others constant. Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$. For our LCG ablations on $S$ and $\lambda$ (Figures 4c and 4d), we set $\alpha=20\%$, $\beta=0.4$ and $\gamma=0.3$ for CIFAR-10, and choose $\alpha=10\%$, $\beta=0.6$ and $\gamma=0.5$ for SVHN, matching the same set of optimal hyperparameters as identified for our selection-based strategies. Figure 5 visualizes the images generated by our guided DDPM fine-tuning techniques, including PCG, LCG-GMM, and LCG-KM, on CIFAR-10 with different numbers of fine-tuning epochs $S\in\{0,5,15,25,35,45\}$. As $S$ is increased, the generated CIFAR-10 images become more difficult to recognize, reflecting a higher degree of uncertainty.
@@ C-A Other Ablations on Hyperparameters
Figures 6a and 6b demonstrate how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary. For CIFAR-10, the per-batch ratio is set to $0.3$, and $\beta$ is fixed at $0.4$. For SVHN, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that as the amount of included unlabeled data increases, the optimal accuracy is reached at a considerably later epoch. When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained in the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
In Figure 6c, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure 6d displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$.
@@
C-B Pre-Trained CLIP as Intermediate Model
In our experiments, we always began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. $f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the intended benefits of the data selection process, primarily to reduce computational and time complexity. In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
To address these issues, we investigate the feasibility of utilizing an alternative pre-trained model in LCS for data selection, specifically by leveraging the CLIP model to extract image embeddings [32]. CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution enables us to select a relevant subset without requiring the training of an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure 7 shows the result when using CLIP instead of the previous intermediate model for unlabeled data selection in SSAT.
TABLE V: Comparison results of SSAT algorithms with varying data selection schemes on SVHN. Instead of pseudo labels, the selected unlabeled data are assigned the ground-truth SVHN labels provided in the dataset.
$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$) #Epochs